OpenStudy (anonymous):

Expand the brackets and collect terms: (X+2)^2 - (X+1)^2 I worked it out to be 1x + 3, but the answer book says 2x + 3, can anyone explain please?

OpenStudy (austinl):

\((x+2)^2=(x+2)(x+2)\) FOIL that out for me.

OpenStudy (anonymous):

FOIL?

OpenStudy (austinl):

\(\color{red}{\text{FOIL}}\) \(\color{red}{\text{F}}\text{irst}\) \(\color{red}{\text{O}}\text{uter}\) \(\color{red}{\text{I}}\text{nner}\) \(\color{red}{\text{L}}\text{ast}\) \((a+b)(c+d)=ac+ad+bc+bd\)

OpenStudy (anonymous):

Ah right, I'm with you, thanks a lot.

OpenStudy (austinl):

No problem :)